# PET-SQL: A Prompt-enhanced Two-stage Text-to-SQL Framework with Cross-consistency

### v3

@@ Abstract
Recent advancements in Text-to-SQL (Text2SQL) emphasize stimulating the large language models (LLM) on in-context learning, achieving significant results. Nevertheless, they face challenges when dealing with verbose database information and complex user intentions. This paper presents a two-stage framework to enhance the performance of current LLM-based natural language to SQL systems. We first introduce a novel prompt representation, called reference-enhanced representation, which includes schema information and randomly sampled cell values from tables to instruct LLMs in generating SQL queries. Then, in the first stage, question-SQL pairs are retrieved as few-shot demonstrations, prompting the LLM to generate a preliminary SQL (PreSQL). After that, the mentioned entities in PreSQL are parsed to conduct schema linking, which can significantly compact the useful information. In the second stage, with the linked schema, we simplify the prompt’s schema information and instruct the LLM to produce the final SQL. Finally, as the post-refinement module, we propose using cross-consistency across different LLMs rather than self-consistency within a particular LLM. Our methods achieve new SOTA results on the Spider benchmark, with an execution accuracy of 87.6%. The codes are released on https://github.com/zhshLii/PETSQL
@@ 1 Introduction
Text-to-SQL (Text2SQL) assists individuals in converting natural language questions (i.e., text) into structure query language (SQL) queries. The task is helpful for question-answering and information-retrieval applications such as traffic and business intelligence [1; 2], as it lowers the requirements for expertise and allows non-professional participants to interact with databases in natural language.
Recent advancements in Text2SQL methods emphasize stimulating the performance of large language models (LLM) on in-context learning with elaborate prompt strategies [3; 4; 5]. Couple with it, typical post-refinements that rely exclusively on executing results or exceptions are proposed, such as self-debugging [1; 6] and self-consistency [4]. However, these approaches encounter difficulties when handling extensive databases across domains and intricate user intentions. First, during the design of the prompt, recent studies focus on the representations/processing of schema information, while lacking a priori attention to cell values format in tables. This relates to the handling of conditionals. For example, the use of conditions ‘sex=‘Male’’, ‘sex=‘male’’, or ‘sex=‘M’’ will produce different results when querying for the gender. In addition, special instructions in the prompt need to be further optimized. Second, to simplify the verbose schema and thus reduce the heavy context, the canonical schema-linking module is usually an independent plugin that instructs LLM to yield question-aware tables and columns for simplification. It consumes the LLM budget once but only produces limited information (i.e., related table and column names). Moreover, the outcomes may lack conciseness, and for coding LLMs like CodeLlama [7], generating non-code (plain) text based on instructions is not their forte. Third, the gains of current post-refinements are not significant. For example, self-debugging is inaccessible if the SQLs generated by a strong LLM have semantic ambiguity rather than syntax errors (our experience with multiple LLMs, especially GPT4). In addition, the investment and production in self-consistency of one LLM are disproportionate, as it calls the same LLM several times (time-consuming and yields many times the cost) at a higher temperature, which does not fundamentally diversify the SQL outputs. The higher temperature may generate more hallucinations, potentially harming the performance of the LLM [8].
In light of the above challenges, in this paper, we present the Prompt-Enhanced Two-stage text-to-SQL framework with cross-consistency (PET-SQL for short), that stimulates the performance of current LLMs without the need to train specific models. Concretely, we first elaborate on a general and effective representation of the prompt, termed reference-enhanced representation. Besides the schema information, the cell values in tables are randomly sampled and added to provide oversight of the value format. We also instruct the LLM to minimize SQL execution time while ensuring correctness. Then, based on the prompt, in the first stage, following the question de-semanticization setting in [9], several question-SQL pairs are retrieved as few-shot demonstrations and prompt LLM to generate the preliminary SQL (PreSQL). After that, the mentioned tables in PreSQL are parsed, which serves as the rationale for schema linking. In the second stage, with the linked schema, we simplify the schema information in the prompt and ask the LLM to yield the final SQL (FinSQL). Finally, we propose to use the cross-consistency of different LLMs rather than the self-consistency of a particular LLM, since we believe that diversity across different LLMs is more beneficial than the repeated calling for an LLM under high temperatures. Overall, the method achieves state-of-the-art in the Spider [10] leaderboard with 87.6% execution accuracy, which achieves a 1% improvement against the DAIL-SQL [4].
The main contributions of the work are summarized as follows:
@@
  • A one-for-all elaborated prompt is designed and its effectiveness is systemically evaluated under various LLMs. Besides the schema information, the cell values in tables and magic instructions are attended to as the prior knowledge in the prompt.
@@ 2 Related Work
Learning-based agents. With the advancement of language models in previous years, multiple methods have been proposed for the Text2SQL task. These early approaches focus on pattern matching between natural language (NL) and SQL statements [11; 12; 13]. Some studies employ relation-aware self-attention mechanisms as encoders to learn representations of questions and schemas. These methods then utilize grammar-based decoders to generate SQL as abstract syntax trees or employ sketch-based decoders to obtain SQL by filling in slots [14; 15]. Subsequently, the advent of pre-trained language models and the fine-tuning paradigm significantly influences these methods [16]. Numerous studies utilize standard sequence-to-sequence models with transformer architectures [17] to translate NL questions into SQL queries in an end-to-end manner [18; 19].
LLM-based in-context learning methods. The development of LLMs has catalyzed substantial transformations in this field. These methods leverage the in-context learning, semantic understanding, and reasoning capabilities of LLMs [20; 21; 22], continuously pushing the boundaries of performance on various evaluation benchmarks, such as Spider [10] and BIRD [23]. For example, C3 [3] achieves a breakthrough by leveraging ChatGPT’s zero-shot learning capability. DIN-SQL  [5] decomposes the task into multiple components, including schema linking, difficulty classification, and SQL generation. This methodology effectively reduces the complexity of decision-making, enabling LLMs to generate more precise SQL statements. DAIL-SQL [4] further enhances the capabilities of LLMs through in-context learning and supervised fine-tuning schemes. In addition, several other studies incorporate advanced reasoning methods, such as chain-of-thought [24] and self-reflexion [25], to enhance the capabilities of LLMs and improve their performance on Text2SQL tasks [26; 27].
@@ 3 Methodology
According to a natural language question $Q$ on a database $D$, the objective of LLM-based Text2SQL is to translate $Q$ into an executable SQL query s{s}. The likelihood of an LLM $\mathcal{M}$ generating a SQL query s{s} can be formally defined as the conditional probability distribution
$\mathbb{P}_{\mathcal{M}}(s|\mathcal{P}({Q},\mathcal{I}(D)))=\prod_{i=1}^{|s|}\mathbb{P}_{\mathcal{M}}(s_{i}|\mathcal{P}(Q,\mathcal{I}(D)),s_{<i}),$ (1)
where $\mathcal{I}(D)$ means the description or information of $D$, such as data definition language. $\mathcal{P}$ is the prompt templete, and $|s|$ is the length of $s$. $s_{i}$ and $s_{<i}$ denotes the $i$-th token and the prefix of $s_{i}$, respectively.
@@
In what follows, we will detail the PET-SQL framework. As shown in Fig. 1, its specifications can be summarized by four points: (1) an elaborated prompt that harnesses the customized instructions, basic database information, and samples in the stored tables (2) instructing the LLM to generate PreSQL, in which some demonstrations are selected from pools using a question similarity-based strategy and then prefixed to the prompt as the few-shot in-context (3) finding question-related tables (schema linking) based on the PreSQL and prompt LLM to yield FinSQL by the linked schema (4) ensuring consistency in the predicted results across multiple LLMs.
Figure 1: The overview of the proposed PET-SQL framework
3.1 Prompting
@@
When instructing LLMs to generate SQL queries, the usage of styles or templates for prompts significantly impacts LLMs’ performance. As [4] recommended, Code Representation ($CR_{p}$) and OpenAI Demonstration ($OD_{p}$) forms mixed with additional information are better choices (shown in Listing 1). To further exploit the potential of LLM, we enrich the prompt based on $OD_{p}$ which is termed reference-enhanced representation ($RE_{p}$).
1OpenAI Demonstration Format:
2### How many singers do we have?
3SELECT count(*) FROM singers
4
5Code Representation Format:
6/* How many singers do we have? */
7SELECT count(*) FROM singers
Listing 1: The simple showcases of OpenAI Demonstration and Code Representation formats.
Building upon $OD_{p}$ which consists of the task instruction, database schema, and question components, $RE_{p}$ performs three modifications: optimization rule (OR), cell value references (CV), and foreign key declarations (FK). Concretely, in the task instruction, we raise a multi-task constraint rule for LLM, that is, highlighting the LLM to “minimize SQL execution time while ensuring correctness” (line 1 in Listing 2). The rationale is not only to focus on execution accuracy but also to make LLM aware of the efficiency of SQL statements. During the generation of efficient SQL statements by LLM, redundant characters and operators can be avoided (which often result in exceptions). Ablation experiments also confirm the effectiveness of the rule. Second, three rows in each table are randomly sampled and inserted in the prompt (lines 7-11 in Listing 2). The sampled cell values as references can help LLM understand the formats and specifications of the database. It relieves the dilemma that being unsure which element to query as the condition, caused by the lack of standardization in filling out different tables. For example, all of the “M”, “Male”, and “male” are feasible representations of gender. Without prior knowledge of cell values, organizing condition statements to query the number of men in a table can be confusing. Thirdly, foreign key relations in the schema are added as suffixes in the prompt (lines 12-15 in Listing 2). It facilitates LLM to recognize the connections between the tables involved in the database, and thus improves the understanding of the user’s intent and automatically selects the appropriate connection in the queries.

### v4

# PET-SQL: A Prompt-Enhanced Two-Round Refinement of Text-to-SQL with Cross-consistency

@@ Abstract
Recent advancements in Text-to-SQL (Text2SQL) emphasize stimulating the large language models (LLM) on in-context learning, achieving significant results. Nevertheless, they face challenges when dealing with verbose database information and complex user intentions. This paper presents a two-round framework to enhance the performance of current LLM-based natural language to SQL systems. We first introduce a novel prompt representation, called reference-enhanced representation, which includes schema information and randomly sampled cell values from tables to instruct LLMs in generating SQL queries. Then, in the first round, question-SQL pairs are retrieved as few-shot demonstrations, prompting the LLM to generate a preliminary SQL (PreSQL). After that, the mentioned entities in PreSQL are parsed to conduct schema linking, which can significantly compact the useful information. In the second round, with the linked schema, we simplify the prompt’s schema information and instruct the LLM to produce the final SQL (FinSQL). Finally, as the post-refinement module, we propose using cross-consistency across different LLMs rather than self-consistency within a particular LLM. Our methods achieve new SOTA results on the Spider benchmark, with an execution accuracy of 87.6%. The codes are released on https://github.com/zhshLii/PETSQL
@@ 1 Introduction
Text-to-SQL (Text2SQL) assists individuals in converting natural language questions (i.e., text) into structure query language (SQL) queries. For example, conditioned on the contents of the user’s question “How many singers do we have?” and the corresponding description for the database schema (e.g., table/column names), the ideal Text2SQL agent can generate SQL statements “SELECT count(*) FROM singer”, which can then be executed on the SQL database engine to retrieve the desired response. The task is helpful for the database question-answering and information-retrieval applications, such as finance, traffic, and business intelligence [1; 2], as it lowers the requirements for expertise and allows non-professional participants to interact with databases in natural language.
Figure 1: (a) A three-stage paradigm of SOTA Text2SQL. (b) We summarize the distribution of error types in DIN-SQL [3]. We elaborate on the methods for each type: PreSQL-based schema linking for errors in schema linking, cell values references-enhanced prompt for incorrect cell values, and cross-consistency for other errors caused by typos or syntax mistakes. (c) Existing schema linking directly lets LLM guess the schema, we let LLM first generate a preliminary SQL and parse table/column entities from PreSQL, increasing accuracy by +4%. (d) Our proposed cross-consistency enables the LLMs to operate at low temperatures, thereby reducing hallucination.
The development trajectory of Text2SQL has gone from rule-based, domain-specific solutions to deep learning-based models, specifically sequence-to-sequence models. Recent large language models (LLM)-based solutions have further pushed the performance to a new high. The recent state-of-the-art high-performers [3; 4] have decomposed the Text2SQL into a series of subtasks. Such a task decomposition has become a new paradigm, and we generalize it as three stages, as shown in Fig. 1(a): (1) pre-processing, including prompt engineering such as zero-shot [5] or few-shot prompting [3], schema linking [3; 6], in-context-learning (ICL) [3], (2) SQL generation, which is based on LLMs such as CodeLlaMa, ChatGPT, GPT-4, (3) post-calibration, including self-consistency [4], self-debugging [1; 7], self-correction [3]. Most of the improvements and designs are happening in the preprocessing and post-calibration, particularly in schema linking (+4% in DIN-SQL with CodeX Davinci) and self-consistency (+0.4% in DAIL-SQL with GPT-4).
However, when handling extensive databases across various domains with different formats and intricate user intentions, there is still large room for improvement in preprocessing and post-calibration.
According to the statistics on the Spider [8], an opensource benchmark database for Text2SQL with 10,181 questions and across 200 complex databases in 138 domains with multiple tables, covering almost all important SQL components including GROUP BY, ORDER BY, HAVING and nested queries, around 60.7% questions are about “WHERE =” or “HAVING =”. However, due to various cell values being formatted, an SQL can be semantically correct but executively wrong. Take a SQL question of querying gender as an example, male can be formatted variously as ‘sex=‘Male’’, ‘sex=‘male’’, or ‘sex=‘M’’, and there are more examples from our experiences, such as 1000 also being ‘1k’ or ‘1000.0’: Such a formatting variance is even more commonplace and critical in real industrial databases, and misaligned formatting in SQL will produce different results or even fail. Thus, in the prompting, we proposed a cell value reference, which provides standardized references in filling out different tables and helps LLM understand the formats and specifications of the database, and in our experiment, it can improve the execution accuracy by maximum +4%.
Furthermore, schema-linking, indicating and narrowing down to which tables and which columns the LLMs should search for the answer, is a critical sub-task and performance-booster for Text2SQL. Although works like DIN-SQL [3] proved that directly using LLMs to infer the tables and columns of interest can work better than an end-to-end trained solution, such as BERT [9], we claim that LLMs even GPT-4, have NOT been trained or supervised fine-tuned with such a domain-specific schema linking task (i.e., input a specific SQL question and output the relevant table names and column names), as it falls outside the mainstream scope of LLMs. As a result, the schema linking still accounts for about 22% of the mistakes in GPT-4, while the conditions in other LLMs are more serious (see Fig. 1(b)). We innovatively proposed a preliminary-SQL (PreSQL)-based schema linking, in which we first generate a rough SQL statement by the LLMs and then parse the mentioned table/column entities as the linking results (see Fig. 1(c)). The rationale behind this is that code (particularly SQL) generation, is a fundamental task across nearly all LLMs, and they may be adequately pre-trained in relevant corpus. Therefore, for LLMs, generating PreSQL is easier and more natural than directly performing schema linking as instructed. Consequently, our PreSQL-based schema linking achieves maximum +4% higher execution accuracy.
Last but not least, two major designs in post-calibration, self-debugging and self-consistency, either have very trivial effects or create more instability in the results, according to our findings. According to our experiences with multiple LLMs, we found that self-debugging is useful for fixing syntax errors, but it tends to be powerless in front of the SQLs generated by strong LLMs, whose problem is semantic ambiguity rather than syntax errors. In terms of self-consistency, the technique that is proposed by the current top-1 opensource solution, is theoretically yielding suboptimal SQL due to its nature: it achieves self-consistency by calling the same LLM several times at a higher temperature and choosing the majority as the final output, which firstly does not fundamentally diversify the SQL outputs, and moreover, the higher temperature generates more hallucinations, potentially yielding a stable but bad output. We instead innovatively propose a “cross-consistency” (in Fig. 1(d)), which calls multiple different LLMs at a low temperature, guaranteeing diversity (due to different LLMs) and further converging to a high-quality result (due to low temperature). We also further proposed two voting schemes, which are (1) native majority voting of the executed results and (2) difficulty-aware voting with distinct candidate LLMs according to the complexity grades of PreSQL. Overall, the method achieves state-of-the-art in the Spider [8] leaderboard with 87.6% execution accuracy, being ranked as the top-1 opensource solution. The main contributions are threefold:
@@
  • An elaborated prompt is designed and its effectiveness is systemically evaluated under various LLMs. Besides the schema information, the cell values in tables and magic instructions are attended to as the prior knowledge in the prompt.
@@ 2 Related Work
Learning-based agents. With the advancement of language models in previous years, multiple methods have been proposed for the Text2SQL task. These early approaches focus on pattern matching between natural language (NL) and SQL statements [10; 11; 12]. Some studies employ relation-aware self-attention mechanisms as encoders to learn representations of questions and schemas. These methods then utilize grammar-based decoders to generate SQL as abstract syntax trees or employ sketch-based decoders to obtain SQL by filling in slots [13; 14]. Subsequently, the advent of pre-trained language models and the fine-tuning paradigm significantly influences these methods [15]. Numerous studies utilize standard sequence-to-sequence models with transformer architectures [16] to translate NL questions into SQL queries in an end-to-end manner [17; 18].
LLM-based in-context learning methods. The development of LLMs has catalyzed substantial transformations in this field. These methods leverage the in-context learning, semantic understanding, and reasoning capabilities of LLMs [19; 20; 21], continuously pushing the boundaries of performance on various evaluation benchmarks, such as Spider [8] and BIRD [22]. For example, C3 [5] achieves a breakthrough by leveraging ChatGPT’s zero-shot learning capability. DIN-SQL  [3] decomposes the task into multiple components, including schema linking, difficulty classification, and SQL generation. This methodology effectively reduces the complexity of decision-making, enabling LLMs to generate more precise SQL statements. DAIL-SQL [4] further enhances the capabilities of LLMs through in-context learning and supervised fine-tuning schemes. In addition, several other studies incorporate advanced reasoning methods, such as chain-of-thought [23] and self-reflexion [24], to enhance the capabilities of LLMs and improve their performance on Text2SQL tasks [25; 26].
@@ 3 Methodology
For a natural language question $Q$ on a database $D$, the objective of LLM-based Text2SQL is to translate $Q$ into an executable SQL query ${s}$. The likelihood of an LLM $\mathcal{M}$ generating a SQL query ${s}$ is defined as a conditional probability distribution, where $\mathcal{I}(D)$ is the description of $D$, $\mathcal{P}$ is the prompt template, and $|s|$ is the length of $s$. $s_{i}$ and $s_{<i}$ are the $i$-th token and the prefix of $s_{i}$:
$\small\mathbb{P}_{\mathcal{M}}(s|\mathcal{P}({Q},\mathcal{I}(D)))=\prod_{i=1}^{|s|}\mathbb{P}_{\mathcal{M}}(s_{i}|\mathcal{P}(Q,\mathcal{I}(D)),s_{<i}),$ (1)
@@
As shown in Fig. 2, our PET-SQL framework mainly includes: (1) an elaborated prompt that harnesses the customized instructions, basic database information, and samples in the stored tables (2) instructing the LLM to generate PreSQL, in which some demonstrations are selected from pools using a question similarity-based strategy and then prefixed to the prompt as the few-shot in-context (3) finding question-related tables (schema linking) based on the PreSQL and prompt LLM to yield FinSQL by the linked schema (4) ensuring consistency in the predicted results across multiple LLMs.
Figure 2: The overview of PET-SQL framework. It first elaborates on an effective SQL-Tailored prompt (Sec. 3.1), which can be enhanced with selected demos as few-shot in-context learning. Then, PreSQL is generated by LLMs for parsing linked schema (Sec. 3.2), where linked entities are retained while unrelated parts are removed from the prompt, thereby simplifying it. After that, the Final SQL can be obtained from LLMs with the concise prompt (Sec. 3.3). Finally, cross-consistency is proposed to post-refine the results generated from multiple LLMs (Sec. 3.4). Our PreSQL-based schema linking shortens the prompt length by 32%, and our cross-consistency boosts the Text2SQL accuracy by 2%.
3.1 Pre-processing: SQL-Tailored Prompting
@@
When instructing LLMs to generate SQL queries, the usage of styles or templates for prompts significantly impacts LLMs’ performance. As [4] recommended, Code Representation ($CR_{p}$) and OpenAI Demonstration ($OD_{p}$) forms mixed with additional information are better choices (shown in Listing 1). To further exploit the potential of LLM, we enrich the prompt based on OpenAI Demonstration ($OD_{p}$), and we name our proposed as Reference-Enhanced representation ($RE_{p}$).
Listing 1: OpenAI Demonstration v.s. Code Representation.
Building upon $OD_{p}$ consisting of the task instruction, database schema, and question components, our proposed $RE_{p}$ performs three modifications: optimization rule (OR), cell value references (CV), and foreign key declarations (FK), which all turn out effective.
Optimization Rule (OR): Concretely, in the task instruction, we raise a multi-task constraint rule for LLM, that is, highlighting the LLM to “minimize SQL execution time while ensuring correctness” (line 1 in Listing 2). The rationale is not only to focus on execution accuracy but also to make LLM aware of the efficiency of SQL statements. During the generation of efficient SQL statements by LLM, redundant characters and operators can be avoided, which often result in exceptions. As our experiment shows, such an optimization rule can boost the performance by as high as +2.3%.
Listing 2: The proposed SQL-Tailored prompt, containing OR, CV, and FK. It is zero-shot without demos.
Cell Value References (CV): As mentioned before, to avoid mistakes from wrong-formatted cell value, three rows in each table are randomly sampled and inserted in the prompt (lines 7-11 in Listing 2). The sampled cell values as references can help LLM understand the database formats and specifications. It relieves the dilemma of being unsure which element to query as the condition caused by the lack of standardization in filling out different tables. Without prior knowledge of cell values, organizing condition statements to query the number of men in a table can be confusing. Such a cell value reference can boost performance by +4.0%.
Foreign Key Declarations (FK): Thirdly, foreign key relations in the schema are added as suffixes in the prompt (lines 12-15 in Listing 2). It facilitates LLM in recognizing the connections between the tables involved in the database, thus improving the understanding of the user’s intent and automatically selecting the appropriate connection in the queries. Such foreign key can boost the performance by as high as +7.5%.